OpenStudy (anonymous):

HELP@&$@#%$!^&!*@ The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 2,500n - 5,250 Part A: What are the zeroes of the above function and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n)? (2 points)

OpenStudy (anonymous):

@jigglypuff314 COULD YOU HELP ? :)

jigglypuff314 (jigglypuff314):

I'm not good at explaining this sort of stuff, sorry :(

OpenStudy (anonymous):

awhh thanks anyways

jigglypuff314 (jigglypuff314):

sorry :( maybe @sourwing could help? :3

OpenStudy (anonymous):

@jigglypuff314 don't be lazy lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

come on @sourwing you can do it ?

jigglypuff314 (jigglypuff314):

:P but I really don't know what the zeros of the function represent of course I know how to find it (with my trusty calculator) but idk what they mean XD

OpenStudy (anonymous):

lol ugh how i hate math could you help on a different question?

OpenStudy (anonymous):

A) P(n) means the profit of selling tickets at n price. So, say P(12) = 10,000. This means the profit is $10,000 if the they decide to sell the concert tickets at $12 each. find zeros means find what value of n gives P(n) = 0. suppose n is, like P(500) = 0, this means there would be no profit if each ticket cost like $500 each. Of course you to find the actual value of n that makes the equations equal 0.

OpenStudy (anonymous):

B and C are just calculator problems.

OpenStudy (anonymous):

i'll leave that to @jigglypuff314 XD.Bye

OpenStudy (anonymous):

ugh ok...

jigglypuff314 (jigglypuff314):

0.o well... P(n) = -250n2 + 2,500n - 5,250 "zeros" would mean -250n^2 + 2500n - 5250 = 0 you can factor out a -250 -250 (n^2 - 10n + 21) = 0 do you know how to factor n^2-10n+21 ???

OpenStudy (anonymous):

yeah

jigglypuff314 (jigglypuff314):

factor it and what do you get? :)

OpenStudy (anonymous):

(n−3)(n−7)

OpenStudy (anonymous):

so what part is that?

jigglypuff314 (jigglypuff314):

yes :) then find -250(n-3)(n-7) = 0 as in what is n-3=0 and n-7=0

jigglypuff314 (jigglypuff314):

erm I think we're still on part a

OpenStudy (anonymous):

is this part b that im currently about to do?

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

n=3 , n=7?

jigglypuff314 (jigglypuff314):

then as sourwig described, if the ticket's price (n) = 3 or 7 in this case, than the profit would be equal to zero

OpenStudy (anonymous):

so is that it?

OpenStudy (anonymous):

for part A.)?

jigglypuff314 (jigglypuff314):

for part a yes :P

OpenStudy (anonymous):

P(n) = -250n2 + 2,500n - 5,250 "zeros" would mean -250n^2 + 2500n - 5250 = 0 you can factor out a -250 -250 (n^2 - 10n + 21) = 0 do you know how to factor n^2-10n+21 would be the steps right

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

erm part b wants you to complete the square so we have -250 (n^2 - 10n + 21) completing the square is like if n^2 - 10n + 25 = (n-5)^2 how would you get n^2 - 10 + 21 like that? ^^

OpenStudy (anonymous):

(n-5)^2 ?

OpenStudy (anonymous):

i dont think i got that right could you help me also can you please show work?

jigglypuff314 (jigglypuff314):

yeah but would you get n^2 - 10n+21 to be like n^2-10n+25 at 4 right? but you can't just add 4 without subtracting it so -250(n^2-10n+21+4-4) -250( (n-5)^2 - 4) -250 (n-5)^2 + 1000

jigglypuff314 (jigglypuff314):

and so n-5=0 solve for n would be your highest profit

OpenStudy (anonymous):

-250 (n-5)^2 + 1000

OpenStudy (anonymous):

do i factor that?

jigglypuff314 (jigglypuff314):

no it would be just like that it would be vertex form: y = a (x-h)^2 + k where h is how much it was shifted right or left and k is how much it was shifted up or down

jigglypuff314 (jigglypuff314):

so n-5=0 solve fo n would be your highest profit

OpenStudy (anonymous):

oh so for part b the highest profit would be n-5=0?

jigglypuff314 (jigglypuff314):

solve for n

OpenStudy (anonymous):

the maximuim profit that is.

jigglypuff314 (jigglypuff314):

but yeah :P

OpenStudy (anonymous):

so n=5 i totally get it

OpenStudy (anonymous):

and wont the axis of symmetry also be 5?

jigglypuff314 (jigglypuff314):

I think so :)

OpenStudy (anonymous):

thanks :) so i have to form a proper equation setence or what not. so for part b i could say so we have -250 (n^2 - 10n + 21) completing the square is like if n^2 - 10n + 25 = (n-5)^2 how would you get n^2 - 10 + 21. so -250(n^2-10n+21+4-4) -250( (n-5)^2 - 4) -250 (n-5)^2 + 1000. so n-5=0 and solving for n you get n=5

jigglypuff314 (jigglypuff314):

yeah :)

OpenStudy (anonymous):

thanks so much could you help me with this equation

jigglypuff314 (jigglypuff314):

erm... 0.o I'm kinda really close to crashing (it's really past my bed time)

OpenStudy (anonymous):

oh ok thanks anyways :) your a great help!

OpenStudy (anonymous):

@ desmonds maybe you should do your own work and not cheat on here...This could be why you're not passing huh?

OpenStudy (anonymous):

@ducky2850 They didn't cheat--they asked for help and went through the steps, instead of just getting the answer. They now know how to do this, which is the whole point of Open Study.